1)An automobile driver increases the velocity at a constant rate from 25 km/h to 55 km/h in 0.50 min. A bicycle rider speeds up at a constant rate from rest to 30 km/h in 0.50 min. Calculate (a) the driver’s acceleration and (b) the rider’s acceleration.
Answer #1

Solution :

For Automobile driver :

Initial velocity (u) = 25 km/h = 6.94 m/s

Final velocity (v) = 55 km/h = 15.28 m/s

Time (t) = 0.50 min = 30 sec

Since : Acceleration (a) = (v - u) / t = {(15.28 m/s) - (6.94 m/s)} / (30 s) = 0.278 m/s2

.

For Bicycle rider :

Initial velocity (u) = 0 m/s

Final velocity (v) = 30 km/h = 8.33 m/s

Time (t) = 0.50 min = 30 sec

Since : Acceleration (a) = (v - u) / t = {(8.33 m/s) - (0 m/s)} / (30 s) = 0.278 m/s2